OpenStudy (anonymous):

In right ∆ABC, m B = 90°, m C = 40°, and BC = 10. What are the other two side lengths of the triangle?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

lets try another picture

OpenStudy (anonymous):



OpenStudy (anonymous):

since this one is a right triangle it is much easier

OpenStudy (anonymous):

we need to label the sides also usually use lower case ones opposite the upper case angles 

OpenStudy (anonymous):

then for example \[\tan(40)=\frac{c}{10}\] so\[c=10\tan(40)\]

OpenStudy (anonymous):

how would you find the answer

OpenStudy (anonymous):

you have to use a calculator

OpenStudy (anonymous):

well i haven't actually used a calculator since about 2004 i would use this http://www.wolframalpha.com/input/?i=10tan%2840%29

OpenStudy (anonymous):

you have lots of choices to find \(b\) you good from there?

OpenStudy (triciaal):

another option is to find b first knowing Sin 90 =1 and using 10/sin 50 = b/sin 90 = c/sin 40 b = 10/sin 50 =13 pythagorean theorem c^2 =13^2 -10^2 =69 so c = rt of 69